\vskip1cm
\begin{thebibliography}{dlPG99}


%\bibitem{Ab} S. Abdulali, Abelian varieties and the general Hodge conjecture, Comp. Math. 109 (1997), 341---355,

\bibitem{An} Y. Andr\'e, Motifs de dimension finie (d'apr\`es S.-I. Kimura, P. O'Sullivan,...), S\'eminaire Bourbaki 2003/2004, Ast\'erisque 299 Exp. No. 929, viii, 115---145,

%\bibitem{And} Y. Andr\'e, Pour une th\'eorie inconditionnelle des motifs, Publications math\'ematiques de l'IHES 83 no 1 (1996), 5---49,

%\bibitem{A} D. Arapura, Motivation for Hodge cycles, Advances in Math. vol. 207 (2006), 762---781,

%\bibitem{Ay} J. Ayoub, Motives and algebraic cycles: a selection of conjectures and open questions, preprint available from user.math.uzh.ch/ayoub/,

%\bibitem{Beau0} A. Beauville, Some remarks on K\"ahler manifolds with $c_1=0$, in: Classification of algebraic and analytic manifolds (Katata, 1982), Birkh\"auser Boston, Boston 1983,

\bibitem{Beau} A. Beauville, Sur l'anneau de Chow d'une vari\'et\'e ab\'elienne, Math. Ann. 273 (1986), 647---651,

%\bibitem{Beau2} A. Beauville, Some surfaces with maximal Picard number, Journal de l'Ecole Polytechnique Tome 1 (2014), 101---116,

\bibitem{Bl2} S. Bloch, Some elementary theorems about algebraic cycles on abelian varieties, Invent. Math. 37 (1976), 215---228,

\bibitem{B} S. Bloch, Lectures on algebraic cycles, Duke Univ. Press Durham 1980,

\bibitem{BKL} S. Bloch, A. Kas and D. Lieberman, Zero cycles on surfaces with $p_g=0$, Comp. Math. 33 no. 2 (1976), 135---145,

\bibitem{BO} S. Bloch and A. Ogus, Gersten's conjecture and the homology of schemes, Ann. Sci. Ecole Norm. Sup. 4 (1974), 181---202,

\bibitem{BS} S. Bloch and V. Srinivas, Remarks on correspondences and algebraic cycles, American Journal of Mathematics Vol. 105, No 5 (1983), 1235---1253,

\bibitem{Bonf} M. Bonfanti, On the cohomology of regular surfaces isogenous to a product of curves with $\chi({\mathcal O}_S)=2$, arXiv:1512.03168v1,

%\bibitem{Bo} C. Borcea, K3 surfaces with involution and mirror pairs of Calabi---Yau manifolds, in: Mirror symmetry,
%II, AMS Studies in Adv. Math. 1, AMS 1997,

\bibitem{Br} M. Brion, Log homogeneous varieties, in: Actas del XVI Coloquio Latinoamericano de Algebra, 
Revista Matem\'atica Iberoamericana, Madrid 2007,
arXiv: math/0609669,

%\bibitem{Bus} N. Buskin, Every rational Hodge isometry between two $K3$ surfaces is algebraic, arXiv:1510.02852v1,

\bibitem{CM} 
M. de Cataldo and L. Migliorini, The Chow groups and the motive of the Hilbert scheme of points on a
surface, Journal of Algebra 251 no. 2 (2002), 824---848,

\bibitem{ChM} F. Charles and E. Markman, The standard conjectures for holomorphic symplectic varieties deformation equivalent to Hilbert schemes of $K3$ surfaces, Comp. Math. 149 (2013),
481---494,

%\bibitem{CD} Clingher and C. Doran, Note on a geometric isogeny of $K3$ surfaces, Int. Math. Research Notices 2011 (2011), 3657---3687,

%\bibitem{CDP} P. Candelas, X. de la Ossa and L. Parkes, Generalized Calabi--Yau manifolds and the mirror of a rigid manifold, Nucl. Phys. B 406 (1993), 115---154,

%\bibitem{CH} A. Corti and M. Hanamura, Motivic decomposition and intersection Chow groups, I, Duke
%Math. J. 103 (2000), 459---522,

%\bibitem{CG} A. Cattaneao and A. Garbagnati, Calabi--Yau $3$--folds of Borcea--Voisin type and elliptic fibrations, arXiv:1312.3481v1,

%\bibitem{CHSW} P. Candelas, G. Horowitz, A. Strominger and E. Witten, Vacuum configurations for superstrings, Nucl. Phys. B 258 (1985), 46---74,

\bibitem{CH} S. Cynk and K. Hulek, Higher--dimensional modular Calabi--Yau manifolds,
Canad. Math. Bull. 50 (4) (2007),  486---503,

\bibitem{D} P. Deligne, La conjecture de Weil pour les surfaces $K3$, Invent. Math. 15 (1972), 206---226,

\bibitem{DM} C. Deninger and J. Murre, Motivic decomposition of abelian schemes and the Fourier transform. J. reine u.
angew. Math. 422 (1991), 201---219,

%\bibitem{Dol} I. Dolgachev, Weighted projective varieties, in: Group actions and vector fields, Vancouver 1981, Springer Lecture Notes in Mathematics 956, Springer Berlin Heidelberg New York 1982,

%\bibitem{FG} S. Filippini and A. Garbagnati, A rigid Calabi--Yau $3$--fold, Adv. Theor. Math. Phys. Vol. 15 No. 6 (2011), 1745---1787,

%\bibitem{LFu} L. Fu, Decomposition of small diagonals and Chow rings of hypersurfaces and Calabi--Yau complete intersections, Advances in Mathematics (2013), 894---924,

%\bibitem{LFu} L. Fu, On the action of symplectic automorphisms on the $CH_0$--groups of some hyper-K\"ahler fourfolds, Math. Z. 280 (2015), 307---334,

%\bibitem{F} W. Fulton, Intersection theory, Springer--Verlag Ergebnisse der Mathematik, Berlin Heidelberg New York Tokyo 1984,

%\bibitem{GG} A. Garbagnati and B. van Geemen, Examples of Calabi--Yau threefolds parametrised by Shimura varieties, Rend. Sem. Mat. Univ. Pol. Torino Vol. 68, 3 (2010), 271---287,

%\bibitem{GG2} A. Garbagnati and B. van Geemen, The Picard---Fuchs equation of a family of Calabi---Yau threefolds without maximal
%unipotent monodromy, International Mathematics Research Notices (2010), 3134---3143,

\bibitem{GaP} A. Garbagnati and M. Penegini, $K3$ surfaces with a non--symplectic automorphism and product--quotient surfaces with cyclic groups, to appear in Rev. Mat. Iberoam.,

%\bibitem{vG} B. van Geemen, Kuga--Satake varieties and the Hodge conjecture, in:
%The Arithmetic and Geometry of Algebraic Cycles, Banff 1998 (B. Gordon et alii, eds.), Kluwer Dordrecht 2000,

%\bibitem{vG2} B. van Geemen, Half twists of Hodge structures of CM--type, J. Math. Soc. Japan Vol. 53 No. 4 (2001), 813---833,

%\bibitem{vGS} B. van Geemen and A. Sarti, Nikulin involutions on $K3$ surfaces, Math. Z. 255 (2007), 731---753,

%\bibitem{G} B. Gordon, Algebraic cycles and the Hodge structure of a Kuga fiber variety, Transactions of the American Math. Society Vol. 336 No 2 (1993), 933---947,

%\bibitem{GS} A. Garbagnati and A. Sarti, Kummer surfaces and $K3$ surfaces with $(\ZZ/2\ZZ)^4$ symplectic action, arXiv:1305.3514, to appear in Rocky Mountain J. of Math.,

\bibitem{GP} V. Guletski\u{\i} and C. Pedrini, The Chow motive of the Godeaux surface, in:
Algebraic Geometry, a volume in memory of Paolo Francia (M.C. Beltrametti,
F. Catanese, C. Ciliberto, A. Lanteri and C. Pedrini, editors),
Walter de Gruyter, Berlin New York, 2002,


%\bibitem{GHM}  B. Gordon, M. Hanamura and J. Murre, Relative Chow-K\"unneth projectors for modular
%varieties, J. Reine u. Angew. Math. 558 (2003), 1---14,

%\bibitem{Hu} D. Huybrechts, Symplectic automorphisms of $K3$ surfaces of arbitrary order, Math. Res. Letters 19 (2012), 947---951,

%\bibitem{WHu} W. Hu, The Chow group of zero cycles for the quotient of certain Calabi--Yau varieties, Houston J. of Math. Vol 38, No 1 (2012), 55---67,

\bibitem{I} F. Ivorra, Finite dimensional motives and applications (following S.-I. Kimura, P. O'Sullivan and others), in:
        Autour des motifs, Asian-French summer school on algebraic geometry and number theory,
       Volume III, Panoramas et synth\`eses, Soci\'et\'e math\'ematique de France 2011,
    
    \bibitem{Iy} J. Iyer, Murre's conjectures and explicit Chow--K\"unneth projectors for varieties with a nef tangent bundle, Transactions of the Amer. Math. Soc. 361 (2008), 1667---1681,

\bibitem{Iy2} J. Iyer, Absolute Chow--K\"unneth decomposition for rational homogeneous bundles and for log homogeneous varieties, Michigan Math. Journal
 Vol.60, 1 (2011), 79---91,
          
%\bibitem{J} U. Jannsen, Mixed motives and algebraic K--theory, Springer Lecture Notes in Mathematics 1400 (1990),

%\bibitem{J} U. Jannsen, 
%Motives, numerical equivalence, and semi-simplicity, Invent. Math. 107(3) (1992), 447---452, 

\bibitem{J2} U. Jannsen, Motivic sheaves and filtrations on Chow groups, in: Motives (U. Jannsen et alii, eds.), Proceedings of Symposia in Pure Mathematics Vol. 55 (1994), Part 1,  

\bibitem{J3} U. Jannsen, Equivalence relations on algebraic cycles, in: The arithmetic and geometry
of algebraic cycles (B. Gordon et alii, eds.), Banff Conference 1998, Kluwer,

\bibitem{J4} U. Jannsen, On finite--dimensional motives and Murre's conjecture, in: Algebraic cycles and motives (J. Nagel and C. Peters, eds.), Cambridge University Press, Cambridge 2007,

%\bibitem{KMP} B. Kahn, J. P. Murre and C. Pedrini, On the transcendental part of the motive of a surface, in: Algebraic cycles and motives (J. Nagel and C. Peters, eds.), Cambridge University Press, Cambridge 2007,

\bibitem{KS} B. Kahn and R. Sebastian, Smash--nilpotent cycles on abelian 3--folds, Math. Res. Letters 16 (2009), 1007---1010,

\bibitem{Kim} S. Kimura, Chow groups are finite dimensional, in some sense,
Math. Ann. 331 (2005), 173---201,

\bibitem{Kim2} S. Kimura, Surjectivity of the cycle map for Chow motives, in: Motives and algebraic cycles (R. de Jeu and J. Lewis, eds.), Amer. Math. Soc. Providence 2009,

\bibitem{KV} S. Kimura and A. Vistoli, Chow rings of infinite symmetric products, Duke Math. J. 85 (1996), 411---430,

\bibitem{K0} S. Kleiman, Algebraic cycles and the Weil conjectures, in: Dix expos\'es sur la cohomologie des sch\'emas, North--Holland Amsterdam, 1968, 359---386, 

\bibitem{K} S. Kleiman, The standard conjectures, in: Motives (U. Jannsen et alii, eds.), Proceedings of Symposia in Pure Mathematics Vol. 55 (1994), Part 1, 

%\bibitem{Ko} K. Koike, Elliptic $K3$ surfaces admitting a Shioda--Inose structure, Comment. Math. Univ. St. Pauli 61 No 1 (2012), 77---86,

%\bibitem{Kun} K. K\"unnemann, A Lefschetz decomposition for Chow motives of abelian schemes, Inv. Math. 113 (1993), 85---102,

%\bibitem{small} R. Laterveer, Algebraic varieties with small Chow groups, J. Math. Kyoto Univ. Vol. 38 No 4 (1998), 673---694,

%\bibitem{yet} R. Laterveer, Yet another version of Mumford's theorem, Archiv Math. (2015),   

%\bibitem{moi2} R. Laterveer, A brief note concerning hard Lefschetz for Chow groups, to appear in Canadian Math. Bulletin, 

%\bibitem{moimult} R. Laterveer, On a multiplicative version of Bloch's conjecture, submitted,

%\bibitem{moi} R. Laterveer, A family of $K3$ surfaces having finite--dimensional motive, to appear in Archiv der Math.,

\bibitem{des} R. Laterveer, Some desultory remarks concerning algebraic cycles and Calabi--Yau threefolds, Rend. Circ. Mat. Palermo 65 (2) (2016), 333---344,

\bibitem{exCY} R. Laterveer, A family of Calabi--Yau threefolds with finite--dimensional motive, submitted to Tokyo Math. J.,

%\bibitem{Li} C. Liedtke, Supersingular $K3$ surfaces are unirational, Invent. Math. 200 (2015), 979---1014,

%\bibitem{Ma} S. Ma, On $K3$ surfaces which dominate Kummer surfaces, Proceedings of the American Math. Soc. 141 No 1 (2013), 131---137,

%\bibitem{Mor} D. Morrison, The Kuga--Satake variety of an abelian surface, J. Algebra 92 (1985), 454---476,

%\bibitem{Mor} D. Morrison, Algebraic cycles on products of surfaces, in: Proceedings, Algebraic Geometry
%Symposium, T\^ohoku University, 1984,

%\bibitem{Mo} D. Morrison, On $K3$ surfaces with large Picard number, Invent. Math. 75 No 1 (1984), 105---121,

%\bibitem{Mor2} D. Morrison, Isogenies between algebraic surfaces with geometric genus one, Tokyo J. Math. Vol 10 no 1 (1987), 179---187,

%\bibitem{Mu} S. Mukai, On the moduli space of bundles on $K3$ surfaces I, in: Vector bundles on algebraic varieties (Bombay, 1984), Tata Institute of Fund. Research, Bombay 1987,

%\bibitem{M} D. Mumford, Rational equivalence of $0$--cycles on surfaces, J. Math. Kyoto Univ. Vol. 9 No 2 (1969), 195---204,

\bibitem{Mur} J. Murre, On a conjectural filtration on the Chow groups of an algebraic variety, parts I and II, Indag. Math. 4 (1993), 177---201,

\bibitem{MNP} J. Murre, J. Nagel and C. Peters, Lectures on the theory of pure motives, Amer. Math. Soc. University Lecture Series 61, Providence 2013,

%\bibitem{Murt} V. Murty, Algebraic cycles on abelian varieties, Duke Math. J. 50 (1983), 487---504,

%\bibitem{NS} J. Nagel and M. Saito, Relative Chow-K\"unneth decompositions for conic bundles and Prym varieties, Int. Math. Res. Not. 2009, no. 16, 2978---3001,

%\bibitem{OG} K. O'Grady, Decomposable cycles and Noether--Lefschetz loci, arXiv:1502.07897v2,

%\bibitem{OS} K. Oguiso and J. Sakurai, Calabi---Yau threefolds of quotient type, Asian J. Math. 5 (2001), 43---78,

%\bibitem{Or} D. Orlov, Equivalences of derived categories and $K3$ surfaces, J. Math. Sci. (New York) 84 (1997), 1361---1381,

%\bibitem{OS} P. O'Sullivan, Algebraic cycles on an abelian variety, J. Reine Angew. Math. 654 (2011), 1---81,

\bibitem{Par} K. Paranjape, Abelian varieties associated to certain K3 surfaces, Comp. Math. 68 (1988), 11---22,

\bibitem{P} C. Pedrini, On the finite dimensionality of a $K3$ surface, Manuscripta Mathematica 138 (2012), 59---72,

%\bibitem{Ped} C. Pedrini, Bloch's conjecture and valences of correspondences for $K3$ surfaces, arXiv:1510.05832v1,

%\bibitem{PP} A. del Padrone and C. Pedrini, Derived categories of coherent sheaves and motives of K3 surfaces, in: Regulators (J. Gil et alii, eds.), Contemp. Math. 571, Amer. Math. Soc., Providence RI, 2012,

\bibitem{PW} C. Pedrini and C. Weibel, Some surfaces of general type for which Bloch's conjecture holds, to appear in: Period Domains, Algebraic Cycles, and Arithmetic, Cambridge Univ. Press, 2015,

%\bibitem{PS} C. Peters and J. Steenbrink, Mixed Hodge structures, Springer--Verlag Ergebnisse der Mathematik, Berlin Heidelberg New York 2008,

%\bibitem{RM} J. Ram\'on Mar\'\i, On the Hodge conjecture for products of certain surfaces, Collect. Math. 59, 1 (2008), 1---26,

%\bibitem{Rito} C. Rito, A note on Todorov surfaces, Osaka Journal of Math. 46(3) (2009), 685---693,

%\bibitem{Roh} C. Rohde, Maximal automorphisms of Calabi---Yau manifolds versus maximally unipotent monodromy,
%Manuscripta Math. 131 (2010), 459---474,

%\bibitem{R} A.A. Rojtman, The torsion of the group of 0--cycles modulo rational equivalence, Annals of Mathematics 111 (1980), 553---569,

\bibitem{Sam} P. Samuel, Relations d'\'equivalence en g\'eom\'etrie alg\'ebrique, in: Proc. Int. Congress Math. 1958, Cambridge Univ. Press New York 1960,

%\bibitem{Schl} U. Schlickewei, The Hodge conjecture for self--products of certain $K3$ surfaces, Journal of Algebra 324 (2010), 507---529,

\bibitem{Sc} T. Scholl, Classical motives, in: Motives (U. Jannsen et alii, eds.), Proceedings of Symposia in Pure Mathematics Vol. 55 (1994), Part 1, 

\bibitem{Seb} R. Sebastian, Smash nilpotent cycles on varieties dominated by products of curves, Comp. Math. 149 (2013), 1511---1518,

\bibitem{Seb2} R. Sebastian, Examples of smash nilpotent cycles on rationally connected varieties, Journal of Algebra 438 (2015), 119---129,

\bibitem{Shi} T. Shioda, The Hodge conjecture for Fermat varieties, Math. Ann. 245 (1979), 175---184,

%\bibitem{SI} T. Shioda and H. Inose, On singular K3 surfaces, in: Complex analysis and algebraic geometry (W. Baily et al., eds.), 
%Iwanami Shoten Tokyo 1977,


\bibitem{Tan} S. Tankeev, On the standard conjecture of Lefschetz type for complex projective threefolds. II, Izvestiya Math. 75:5 (2011), 1047---1062,

\bibitem{V} C. Vial, Algebraic cycles and fibrations, Documenta Math. 18 (2013), 1521---1553,

\bibitem{V2} C. Vial, Projectors on the intermediate algebraic Jacobians, New York J. Math. 19 (2013), 793---822,

\bibitem{V3} C. Vial, Remarks on motives of abelian type, to appear in Tohoku Math. J.,

\bibitem{V4} C. Vial, Niveau and coniveau filtrations on cohomology groups and Chow groups, Proceedings of the LMS 106(2) (2013), 410---444,

\bibitem{V5} C. Vial, Chow--K\"unneth decomposition for $3$-- and $4$--folds fibred by varieties with trivial Chow group of zero--cycles, J. Alg. Geom. 24 (2015), 51---80,

\bibitem{Voe} V. Voevodsky, A nilpotence theorem for cycles algebraically equivalent to zero, Internat. Math. Research Notices 4 (1995), 187---198,

\bibitem{V9} C. Voisin, Remarks on zero--cycles of self--products of varieties, in: Moduli of vector bundles, Proceedings of the Taniguchi Congress  (M. Maruyama,  ed.), Marcel Dekker New York Basel Hong Kong 1994,

%\bibitem{V10} C. Voisin, Remarks on filtrations on Chow groups and the Bloch conjecture,
%Annali di matematica pura ed applicata 183 (2004), 421---438,

%\bibitem{V20} C. Voisin, Miroirs et involutions sur les surfaces K3, Ast\'erisque 218 (1993), 273---323,

%\bibitem{V12} C. Voisin, Sur les z\'ero--cycles de certaines hypersurfaces munies d'un automorphisme, Annali della Scuola Norm. Sup. di Pisa Vol. 29 (1993), 473---492,

%\bibitem{V13} C. Voisin, Chow rings and decomposition theorems for $K3$ surfaces and Calabi---Yau hypersurfaces, Geom. Topol.
%16 (2012), 433---473,

%\bibitem{V11} C. Voisin, Symplectic involutions of $K3$ surfaces act trivially on $CH_0$, Documenta Math. 17 (2012), 851---860,

%\bibitem{V0} C. Voisin, The generalized Hodge and Bloch conjectures are equivalent for general complete intersections, Ann. Sci. Ecole Norm. Sup. 46, fascicule 3 (2013), 449---475,

\bibitem{V8} C. Voisin, Bloch's conjecture for Catanese and Barlow surfaces, J. Differential Geometry 97 (2014), 149---175,

\bibitem{Vo} C. Voisin, Chow Rings, Decomposition of the Diagonal, and the Topology of Families, Princeton University Press, Princeton and Oxford, 2014,

\bibitem{V1} C. Voisin, The generalized Hodge and Bloch conjectures are equivalent for general complete intersections, II, J. Math. Sci. Univ. Tokyo  22 (2015), 491---517,

\bibitem{Xu} Z. Xu, Algebraic cycles on a generalized Kummer variety, arXiv:1506.04297v1.

\end{thebibliography}

\documentclass[a4paper,12pt,final]{amsart}
\usepackage{times,a4wide,mathrsfs,amssymb}

\newcommand{\C}{\mathbb{C}}
\newcommand{\ZZ}{\mathbb{Z}}

\newcommand{\QQ}{\mathbb{Q}}
\newcommand{\NN}{\mathbb{N}}

\newcommand{\MM}{\mathcal M}

\newcommand{\wt}{\widetilde}
\newcommand{\ima}{\hbox{Im}}
\newcommand{\rom}{\romannumeral}

\newcommand\undermat[2]{
  \makebox[0pt][l]{$\smash{\underbrace{\phantom{
    \begin{matrix}#2\end{matrix}}}_{\text{$#1$}}}$}#2}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{conjecture}[theorem]{Conjecture}
\newtheorem{remark}[theorem]{Remark}
\newtheorem{definition}[theorem]{Definition}
\newtheorem{convention}{Conventions}

\newtheorem{notation}[theorem]{Notation}

\newtheorem{nonumbering}{Theorem}

\newtheorem{nonumberingt}{Acknowledgements}

\begin{document}
\author[Robert Laterveer]
{Robert Laterveer}

\address{Institut de Recherche Math\'ematique Avanc\'ee,
CNRS -- Universit\'e 
de Strasbourg,\
7 Rue Ren\'e Des\-car\-tes, 67084 Strasbourg CEDEX,
FRANCE.}
\email{robert.laterveer@math.unistra.fr}

\title{Some new examples of smash--nilpotent algebraic cycles}

\begin{abstract} Voevodsky has conjectured that numerical equivalence and smash--equivalence coincide for algebraic cycles on any smooth projective variety. Building on work of Vial and Kahn--Sebastian, we give some new examples of varieties where Voevodsky's conjecture is verified.
\end{abstract}

\keywords{Algebraic cycles, Chow groups, motives, finite--dimensional motives, smash--nilpotence, abelian varieties}

\subjclass{Primary 14C15, 14C25, 14C30. Secondary 14J30, 14J32, 14K99}

\maketitle

\section{Introduction}

Let $X$ be a smooth projective variety over $\C$. There exist numerous adequate equivalence relations (in the sense of \cite{Sam}) on the group of algebraic cycles on $X$, ranging from rational equivalence (the finest) to numerical equivalence (the coarsest). Rational equivalence gives rise to the Chow groups $A^j(X):=CH^j(X)_{\QQ}$  (i.e., codimension $j$ cycles with rational coefficients modulo rational equivalence). The other equivalence relations give rise to subgroups $A^j_\sim$, for example there are subgroups
  \[   A^j_{alg}(X)\subset A^j_\otimes(X)\subset A^j_{hom}(X)\subset A^j_{num}(X)\]
  of cycles algebraically resp. smash--nilpotent resp. homologically resp. numerically trivial.
Here, the first inclusion is a theorem of Voevodsky \cite{Voe} and Voisin \cite{V9}, and the last inclusion is the subject of one of the standard conjectures \cite{K}.
More ambitiously, Voevodsky has conjectured that $A^j_\otimes(X)$ and $A^j_{num}(X)$ should coincide \cite{Voe}.

Not a great deal is known about this conjecture of Voevodsky's; most results focus on $1$--cycles. For instance, Voevodsky's conjecture has been proven for $1$--cycles on varieties rationally dominated by products of curves \cite{Seb}, \cite[Proposition 2]{Seb2} (this is further generalized by \cite[Theorem 3.17]{V3}).

In this note (which is inspired by \cite{Seb}, \cite{Seb2} and particularly \cite{V3}), we aim for results for cycles in other dimensions by restricting attention to very special varieties.
 The main result is as follows:

\begin{nonumbering}[$\approx$ theorem \ref{main}\footnote{(The actual statement of theorem \ref{main} is somewhat more general, but this simplified version suffices for many applications.)}] Let  
 $X$ be a smooth projective variety.
 Assume that $X$ is dominated by a product of curves, and that the even cohomology of $X$ verifies
     \[ H^{2i}(X,\QQ)=\wt{N}^{i-1} H^{2i}(X,\QQ)\ .\]
      Then
    \[ A^j_\otimes(X)=A^j_{num}(X)\ \ \hbox{for\ all\ }j\ .\]
      \end{nonumbering}

Here $\wt{N}^\ast$ denotes Vial's niveau filtration, which is a variant of the coniveau filtration (cf. \cite{V4} and section \ref{niv} below). Conjecturally, the condition $H^{2i}(X,\QQ)=\wt{N}^{i-1}$ is equivalent to having $H^{2i}(X,\C)=F^{i-1} H^{2i}(X,\C)$, where $F^\ast$ is the Hodge filtration.

 Examples of varieties to which theorem \ref{main} applies include the following: Fermat hypersurfaces of odd dimension; products of type $X^2_d\times X^n_{d^\prime}$ with $n$ odd  (where $X^n_d$ denotes a Fermat hypersurface of dimension $n$ and degree $d$). 
  Some more examples where theorem \ref{main} applies are given in corollary \ref{ex}.



\vskip0.6cm

\begin{convention} In this note, the word {\sl variety\/} will refer to a reduced irreducible scheme of finite type over $\C$. A {\sl subvariety\/} is a (possibly reducible) reduced subscheme which is equidimensional. 

{\bf All Chow groups will be with rational coefficients}: we denote by $A_jX$ the Chow group of $j$--dimensional cycles on $X$ with $\QQ$--coefficients; for $X$ smooth of dimension $n$ the notations $A_jX$ and $A^{n-j}X$ will be used interchangeably. 

The notations $A^j_{hom}(X)$, $A^j_{num}(X)$, $A^j_{AJ}(X)$, $A^j_{alg}(X)$ and $A^j_\otimes(X)$ will be used to indicate the subgroups of homologically trivial, resp. numerically trivial, resp. Abel--Jacobi trivial resp. algebraically trivial, resp. smash--nilpotent cycles.
The contravariant category of Chow motives (i.e., pure motives with respect to rational equivalence as in \cite{Sc}, \cite{MNP}) will be denoted $\MM_{\rm rat}$.



We will write $H^j(X)$ 
and $H_j(X)$ 
to indicate singular cohomology $H^j(X,\QQ)$,
resp. Borel--Moore homology $H_j(X,\QQ)$.

\end{convention}

\section{Preliminary}

\subsection{Motives of abelian type}

We refer to \cite{Kim}, \cite{An}, \cite{I}, \cite{J4}, \cite{MNP} for the definition of finite--dimensional motive. 
An essential property of varieties with finite--dimensional motive is embodied by the nilpotence theorem:

\begin{theorem}[Kimura \cite{Kim}]\label{nilp} Let $X$ be a smooth projective variety of dimension $n$ with finite--dimensional motive. Let $\Gamma\in A^n(X\times X)_{}$ be a correspondence which is numerically trivial. Then there is $N\in\NN$ such that
     \[ \Gamma^{\circ N}=0\ \ \ \ \hbox{\rm in}\ \  A^n(X\times X)_{}\ \]
     (here, $\circ$ indicates composition of correspondences).
\end{theorem}

 Actually, the nilpotence property (for all powers of $X$) could serve as an alternative definition of finite--dimensional motive, as shown by Jannsen \cite[Corollary 3.9]{J4}.

\begin{conjecture}[Kimura \cite{Kim}]\label{findim} Every smooth projective variety has finite--dimensional motive.
\end{conjecture}

We are still far from knowing this, but at least there are quite a few non--trivial examples:
 
\begin{remark}\label{exfindim} 
The following varieties have finite--dimensional motive: abelian varieties, varieties dominated by products of curves \cite{Kim}, $K3$ surfaces with Picard number $19$ or $20$ \cite{P}, surfaces not of general type with vanishing geometric genus \cite[Theorem 2.11]{GP}, Godeaux surfaces \cite{GP}, certain surfaces of general type with $p_g=0$ \cite{PW}, \cite{V8}, Hilbert schemes of surfaces known to have finite--dimensional motive \cite{CM}, generalized Kummer varieties \cite[Remark 2.9(\rom2)]{Xu},
 3--folds with nef tangent bundle \cite{Iy} (an alternative proof is given in \cite[Example 3.16]{V3}), 4--folds with nef tangent bundle \cite{Iy2}, log--homogeneous varieties in the sense of \cite{Br} (this follows from \cite[Theorem 4.4]{Iy2}), certain 3--folds of general type \cite[Section 8]{V5}, varieties of dimension $\le 3$ rationally dominated by products of curves \cite[Example 3.15]{V3}, varieties $X$ with $A^i_{AJ}(X)_{}=0$ for all $i$ \cite[Theorem 4]{V2}, products of varieties with finite--dimensional motive \cite{Kim}.
\end{remark}

\begin{definition}\label{abtype} Let $X$ be a smooth projective variety of dimension $n$. We say that $X$ has {\em motive of abelian type\/} if $h(X)\in\MM_{\rm rat}$ is in the subcategory generated by the motives of curves.

\end{definition}

\begin{remark} It follows from the fact that curves have finite--dimensional motive that ``motive of abelian type'' implies ``finite--dimensional motive''. The converse is probably not true
(many motives are {\em not\/} of abelian type, cf. \cite[7.6]{D}), yet it is a (somewhat embarassing) fact that all known finite--dimensional motives happen to be of abelian type.


\end{remark}

Various characterizations of motives of abelian type are given in \cite{V3}. One of these is as follows:

\begin{proposition}[Vial \cite{V3}]\label{v1} Let $X$ be a smooth projective variety of dimension $n$. The motive of $X$ is of abelian type if and only if $A^j_{alg}(X)$ is generated, via correspondences, by Chow groups of products of curves, for all $j>\lceil {n\over 2}\rceil$.
\end{proposition}

\begin{proof} This follows from \cite[Theorem 5]{V3}.
\end{proof}

\begin{proposition}[Vial \cite{V3}]\label{v2} Let $X$ be a smooth projective variety of dimension $n$, and assume $X$ has motive of abelian type.  
  Then the motive of $X$ is isomorphic to a direct summand
  \[  h(X)\ \subset\ \bigoplus_j h(A_j)(m_j)\ \ \ \hbox{\rm in}\ \MM_{\rm rat}\ ,\]
  where the $A_j$ are abelian varieties.
  \end{proposition} 

\begin{proof} It suffices to note that for motives of abelian type there is an inclusion
   \[ h(X)\ \subset\ \bigoplus_j h(M_j)(m_j) \ \ \ \hbox{in}\ \MM_{\rm rat}\ ,\]
   where $M_j$ is a product of curves $C_1\times\cdots\times C_{r_j}$ (this follows from \cite[Theorem 4]{V3}, plus \cite[Theorem 3.11]{V3} applied with $l=d:=\dim X$).
   It is well--known this implies proposition \ref{v2}. 
  
  (Indeed, for some $n_i\ge 2g(C_i)$ let $C_i^{[n_i]}$ denote the $n_i$--th symmetric product, and let $J_i$ denote the Jacobian of $C_i$. There exist morphisms
  \[   M:=C_1\times \cdots C_r\ \to\ C_1^{[n_1]}\times\cdots\times C_r^{[n_r]}\ \to\ J_1\times\cdots \times J_r\ .\]
  The first arrow identifies $h(M)$ with a direct summand of $h(C_1^{[n_1]}\times\cdots\times C_r^{[n_r]})$ \cite{KV}. The second arrow is a composition of projective bundles, so the motive $h(C_1^{[n_1]}\times\cdots\times C_r^{[n_r]})$ identifies with a sum of shifted motives of $J_1\times\cdots\times J_r$.)
   \end{proof}

\subsection{Lefschetz standard conjecture}

\begin{notation}
Let $X$ be a smooth projective variety of dimension $n$, and $h\in H^2(X,\QQ)$ the class of an ample line bundle. The hard Lefschetz theorem asserts that the map
  \[  L^{n-i}\colon H^i(X)\to H^{2n-i}(X)\]
  obtained by cupping with $h^{n-i}$ is an isomorphism, for any $i< n$.
  \end{notation}
  
   One of the standard conjectures asserts that the inverse isomorphism is algebraic:

\begin{definition} Given a variety $X$, we say that $B(X)$ holds if for all ample $h$, and all $i<n$ the isomorphism 
  \[  (L^{n-i})^{-1}\colon 
  H^{2n-i}(X)\stackrel{\cong}{\rightarrow} H^i(X)\]
  is induced by a correspondence.
 \end{definition}  
 
 \begin{remark} It is known that $B(X)$ holds for the following varieties: curves, surfaces, abelian varieties \cite{K0}, \cite{K}, threefolds not of general type \cite{Tan}, hyperk\"ahler varieties of 
 $K3^{[n]}$--type \cite{ChM}, $n$--dimensional varieties $X$ which have $A_i(X)_{}$ supported on a subvariety of dimension $i+2$ for all $i\le{n-3\over 2}$ \cite[Theorem 7.1]{V}, $n$--dimensional varieties $X$ which have $H_i(X)=N^{\llcorner {i\over 2}\lrcorner}H_i(X)$ for all $i>n$ \cite[Theorem 4.2]{V2}, products and hyperplane sections of any of these \cite{K0}, \cite{K}.
 \end{remark}
 
 \begin{remark}\label{B}
 Let $X$ be a variety with motive of abelian type. Then $B(X)$ holds. This is because the standard conjecture $B$ can also be formulated for motives. Since $B(A)$ holds for abelian varieties, it also holds for direct summands of a sum of twisted motives of abelian varieties, hence for varieties with motive of abelian type.
 It follows that the standard conjectures $C(X)$ (i.e., algebraicity of the K\"unneth components) and $D(X)$ (i.e., homological and numerical equivalence coincide on $X$ and on 
 $X\times X$) also hold \cite{K0}, \cite{K}. 
 \end{remark}

\subsection{Niveau filtration}\label{niv}

\begin{definition}[Coniveau filtration \cite{BO}]\label{con} Let $X$ be a quasi--projective variety. The {\em coniveau filtration\/} on cohomology and on homology is defined as
  \[\begin{split}   N^c H^i(X,\QQ)&= \sum \ima\bigl( H^i_Y(X,\QQ)\to H^i(X,\QQ)\bigr)\ ;\\
                           N^c H_i(X,\QQ)&=\sum \ima \bigl( H_i(Z,\QQ)\to H_i(X,\QQ)\bigr)\ ,\\
                           \end{split}\]
   where $Y$ runs over codimension $\ge c$ subvarieties of $X$, and $Z$ over dimension $\le i-c$ subvarieties.
 \end{definition}

Vial introduced the following variant of the coniveau filtration:

\begin{definition}[Niveau filtration \cite{V4}] Let $X$ be a smooth projective variety. The {\em niveau filtration} on homology is defined as
  \[ \wt{N}^j H_i(X)=\sum_{\Gamma\in A_{i-j}(Z\times X)_{}} \ima\bigl( H_{i-2j}(Z)\to H_i(X)\bigr)\ ,\]
  where the union runs over all smooth projective varieties $Z$ of dimension $i-2j$, and all correspondences $\Gamma\in A_{i-j}(Z\times X)_{}$.
  The niveau filtration on cohomology is defined as
  \[   \wt{N}^c H^iX:=   \wt{N}^{c-i+n} H_{2n-i}X\ .\]
  
\end{definition}

\begin{remark}\label{is}
The niveau filtration is included in the coniveau filtration:
  \[ \wt{N}^j H^i(X)\subset N^j H^i(X)\ .\] 
  These two filtrations are expected to coincide; indeed, Vial shows this is true if and only if the Lefschetz standard conjecture is true for all varieties \cite[Proposition 1.1]{V4}. 
  
  Using the truth of the Lefschetz standard conjecture in degree $\le 1$, it can be checked \cite[page 415 "Properties"]{V4} that the two filtrations coincide in a certain range:
  \[  \wt{N}^j H^i(X)= N^j H^iX\ \ \ \hbox{for\ all\ }j\ge {i-1\over 2} \ .\]
  \end{remark}

 \begin{lemma}\label{split} Let $X$ be a smooth projective variety of dimension $n$ such that $B(X)$ holds. Suppose 
  \[ H^{2i}(X)= \wt{N}^{ i-1} H^{2i}(X) \]
  for some $i$. Then there exists a smooth projective surface $S_{}$ and correspondences $\Gamma_{2i}\in A^{n+1-i}(X\times S_{})$, $\Psi_{2i}\in A^{i+1}(S_{}\times X)$ such that
   \[ \pi_{2i}=\Psi_{2i}\circ \Gamma_{2i}\ \ \ \hbox{in}\ H^{2n}(X\times X)\ .\]
   \end{lemma}
   
   \begin{proof} This follows readily from the arguments contained in \cite{V4}. Indeed, by assumption there exists a surface $S_{}$ and a correspondence $\Psi_{2i}\in A^{i+1}(S_{}\times X)$
   such that
     \[ H^{2i}(X)=(\Psi_{2i})_\ast H^2(S_{})\ .\]
     This means that the homomorphism of motives
     \[  \Psi_{2i}\colon\ \ (S_{},\pi_2,0)\ \to\  (X,\pi_{2i},0)\ \ \ \hbox{in}\ \MM_{\rm hom} \]
     is surjective (i.e., 
     \[ (\Psi_{2i}\times \Delta_M)_\ast\colon H^\ast(S_{}\times M)\ \to\  (\pi_{2i}\times\Delta_M)_\ast H^\ast(X\times M)\]
      is surjective for all smooth projective varieties $M$). On the other hand, the motives $(S_{},\pi_2,0)$ and $ (X,\pi_{2i},0)$ lie in a subcategory $\MM_{\rm hom}^\circ\subset \MM_{\rm hom}$ which is semi--simple (one can define $\MM_{\rm hom}^\circ$ as the smallest full subcategory containing the motives of all varieties $M$ for which $B(M)$ is known). As such, there is a left--inverse to $\Psi_{2i}$; this gives the correspondence $\Gamma_{2i}$ with the property that $\Psi_{2i}\circ \Gamma_{2i}=\pi_{2i}$.
        \end{proof}

\subsection{Smash--nilpotence}

\begin{definition} Let $X$ be a smooth projective variety. A cycle $a\in A^r(X)$ is called {\em smash--nilpotent\/} 
if there exists $m\in\NN$ such that
  \[ \begin{array}[c]{ccc}  a^m:= &\undermat{(m\hbox{ times})}{a\times\cdots\times a}&=0\ \ \hbox{in}\  A^{mr}(X\times\cdots\times X)_{}\ .
  \end{array}\]
  \vskip0.6cm

We will write $A^r_\otimes(X)\subset A^r(X)$ for the subgroup of smash--nilpotent cycles.
\end{definition}

\begin{conjecture}[Voevodsky \cite{Voe}]\label{voe} Let $X$ be a smooth projective variety. Then
  \[  A^r_{num}(X)\ \subset\ A^r_\otimes(X)\ \ \ \hbox{for\ all\ }r\ .\]
  \end{conjecture}

\begin{remark} It is known \cite[Th\'eor\`eme 3.33]{An} that conjecture \ref{voe} implies (and is strictly stronger than) conjecture \ref{findim}.
\end{remark}

The most general result concerning smash--nilpotence is the following:

\begin{theorem}[Voevodsky \cite{Voe}, Voisin \cite{V9}]\label{VV} Let $X$ be a smooth projective variety. Then
  \[ A^r_{alg}(X)\ \subset\ A^r_\otimes(X)\ \ \ \hbox{for\ all\ }r\ .\]
  \end{theorem}
  
  In particular, it follows from theorem \ref{VV} that conjecture \ref{voe} is true for $r=1$ and for $r=\dim X$. Another useful result is the following (this is \cite[Proposition 1]{KS}, which builds on results of Kimura's \cite{Kim}):

\begin{theorem}[Kahn--Sebastian \cite{KS}]\label{skew} Let $A$ be an abelian variety. Assume $a\in A^r(A)$ is {\em skew\/}, i.e. $(-1)^\ast(a)=-a$ in $A^r(A)$. Then $a\in A^r_\otimes(A)$.
\end{theorem}

\section{Main result}

This section contains the proof of our main result (stated in somewhat more general form than in the introduction):

\begin{theorem}\label{main} Let $X$ be a smooth projective variety of dimension $n$. Assume


\noindent
{(\rom1)} $X$ has motive of abelian type;


\noindent
{(\rom2)} $H^{2i}(X)=\wt{N}^{i-1} H^{2i}(X)$ for all $i\le n/2$.

Then Voevodsky's conjecture is true for $X$, i.e.
  \[ A^r_\otimes(X)=A^r_{num}(X)\ \ \hbox{for\ all\ }r\ .\]
  \end{theorem}

\begin{proof} 
Let us denote
  \[ Z^r(X):= { A^r_{num}(X)\over A^r_\otimes(X) }\ .\]

By assumption (\rom1), the K\"unneth components $\pi_i$ of $X$ are algebraic (remark \ref{B}). 
By assumption (\rom2) and lemma \ref{split}, any ``even'' K\"unneth component $\pi_{2i}$ with $i\le n/2$
factors over a surface, i.e. there exists a surface $S_{2i}$ and correspondences $\Gamma_{2i}\in A^{n+1-i}(X\times S_{2i})$, $\Psi_{2i}\in A^{i+1}(S_{2i}\times X)$ such that
   \[ \pi_{2i}=\Psi_{2i}\circ \Gamma_{2i}\ \ \ \hbox{in}\ H^{2n}(X\times X)\ .\]
  
  We now lift the $\pi_i$ to the level of rational equivalence in the following way: for the even components we choose 
    \[    \Pi_{2i}:=\begin{cases} \ \ \  \Psi_{2i}\circ \Gamma_{2i}  \ \ \ \ \ \ \ \ \ \hbox{in}\ A^{n}(X\times X) &\hbox{if\ }i\le n/2\ ;\\
                                                   {}^t  (\Psi_{2n-2i}\circ \Gamma_{2n-2i})  \ \ \ \hbox{in}\ A^{n}(X\times X)&\hbox{if}\ i> n/2\ , \\
                                                   \end{cases}\]
    Here $\Psi_{2i}, \Gamma_{2i}$ are correspondences to and from a surface $S_{2i}$ as above, and ${}^t()$ denotes the transpose of a correspondence.
   For the odd K\"unneth components $\pi_{2i+1}$, we take arbitrary lifts $\Pi_{2i+1}\in A^n(X\times X)$ of the $\pi_{2i+1}$, subject only to the condition that
   \[  \Delta_X={\displaystyle \sum_{i=0}^{2n}} \Pi_i\ \ \hbox{in}\ A^n(X\times X) \]
   (i.e., we define the last $\Pi_{2i+1}$ as a difference of cycle classes). Note that our $\Pi_i\in A^n(X\times X)$ need {\em not\/} be idempotents.
  
  We now remark that
    \[ (\Pi_{2i})_\ast\colon\ \ Z^r(X)\ \to\ Z^r(X) \]
    factors over $Z^\ast(S_{2i})$, which is $0$ since $S_{2i}$ is a surface, and so
    \[  (\Pi_{2i})_\ast=0\colon\ \ Z^r(X)\ \to\ Z^r(X)\ \ \ \hbox{for\ all\ }i\ \hbox{and\ all\ }r\ .\]
   It follows that
    \[ (\Delta_X)_\ast=({\displaystyle \sum_{i\ {\rm odd}}} \Pi_i)_\ast\colon\ \ Z^r(X)\ \to\ Z^r(X)\ .\]
    For later use, let us note that this last equality also implies 
    \begin{equation}\label{power}
    ({\displaystyle \sum_{i\ {\rm odd}}} \Pi_i)_\ast=\Bigl(({\displaystyle \sum_{i\ {\rm odd}}} \Pi_i)^{\circ m}\Bigr){}_\ast=\hbox{id}\colon\ \ Z^r(X)\ \to\ Z^r(X)\ ,\ \ \hbox{for\ all\ }m\in \NN\ .
    \end{equation}

Assumption (\rom1) implies the motive of $X$ identifies with a direct summand
  \[ h(X)\ \subset\ \bigoplus_{j=1}^s h(A_j)(m_j)\ \ \ \hbox{in}\ \ \MM_{\rm rat}\ ,\]
  where the $A_j$ are abelian varieties (proposition \ref{v2}). This formally implies that there exist correspondences
    \[  \begin{split} \Gamma_1&={\displaystyle\sum_j} \Gamma_1^j\ \ \in \bigoplus_j A^\ast(X\times A_j)\ ,\\
                            \Gamma_2&={\displaystyle\sum_j} \Gamma_2^j\ \ \in \bigoplus_j A^\ast(A_j\times X)\\
                            \end{split}\]
                     such that
            \[ \Gamma_2\circ \Gamma_1={\displaystyle\sum_j} \Gamma_2^j\circ \Gamma_1^j=\Delta_X\ \ \hbox{in}\ A^n(X\times X)\ .\]
       In particular, for any $i$ we also have that the composition
       \[  H^{2i+1}(X)\ \xrightarrow{(\Gamma_1)_\ast}\ \bigoplus_j H^{2i+1+2c_j}(A_j)\ \xrightarrow{(\Gamma_2)_\ast}\ H^{2i+1}(X)\]
       is equal to the identity (here $c_j$ is some integer, dependent on $n$ and $\dim A_j$ and $m_j$).     
       But this composition is the same as
       \[    \begin{split}  H^{2i+1}(X)\ \xrightarrow{(\Gamma_1)_\ast}\ \bigoplus_j H^{2i+1+2c_j}(A_j)&\ 
       \xrightarrow{\bigl((\pi_{2i+1+2c_1}^{A_1})_\ast,\ldots, (\pi_{2i+1+2c_s}^{A_s})_\ast\bigr)}\ \\
           &\bigoplus_j H^{2i+1+2c_j}(A_j)\ \xrightarrow{(\Gamma_2)_\ast}\ H^{2i+1}(X)\ ,\\
           \end{split}\]
        where the $\pi^{A_j}_i$ denote the Chow--K\"unneth decomposition of \cite{Mur} for the abelian variety $A_j$.  
           
  That is, we have a homological equivalence
    \[  \Pi_{2i+1}=\Gamma_2\circ \Gamma_1\circ \Pi_{2i+1}=
    {\displaystyle\sum_j}\ \Gamma_2^j\circ \pi^A_{2i+1+2c_j}\circ \Gamma_1^j\circ \Pi_{2i+1}\ \ \hbox{in}\ H^{2n}(X\times X)\ .\]
   Taking the sum over all odd K\"unneth components, we find that
     \[ {\displaystyle  \sum_{i\ {\rm odd}}} \Pi_{i}     - {\displaystyle  \sum_{i\ {\rm odd}}}  {\displaystyle \sum_j} \Gamma_2^j\circ \pi^{A_j}_{2i+1+2c_j}\circ \Gamma_1^j\circ \Pi_{2i+1}\ \ \in\  A^n_{}(X\times X)\ \]
     is homologically trivial. But then (since $X$ has finite--dimensional motive), it follows from theorem \ref{nilp} this cycle is nilpotent: there exists $N\in\NN$ such that
     \[ \Bigl({\displaystyle  \sum_{i\ {\rm odd}}} \Pi_{i}     -{\displaystyle \sum_{i\ {\rm odd}}}{\displaystyle \sum_j} \Gamma_2^j\circ \pi^{A_j}_{2i+1+2c_j}\circ \Gamma_1^j\circ \Pi_{2i+1}\Bigr)^{\circ N}=0\ \ \in\  A^n_{}(X\times X)\ .\]
          Developing this expression, we obtain
      \[     \bigl( {\displaystyle \sum_{i\ {\rm odd}}}\Pi_{i}\bigr)^{\circ N}=  Q_1+\cdots + Q_{N^\prime}
       \ \ \hbox{in}\ A^n(X\times X)\ ,\]
       where each $Q_s$ is a composition of correspondences in which some $\pi^{A_j}_{2i+1+2c_j}$ occurs at least once, i.e.
       \[ Q_s= (\hbox{something})\circ \pi^{A_j}_{\ell}   \circ (\hbox{something})\ \ \ \hbox{in}\ A^n(X\times X)\ \ \ ,\hbox{with\ $\ell$\ odd}.\] 
                This equality implies in particular that both sides act in the same way on $Z^r(X)$ for any $r$, i.e.
      \[ \Bigl( \bigl( {\displaystyle \sum_{i\ {\rm odd}}}\Pi_{i}\bigr)^{\circ N}\Bigr){}_\ast= ( {\displaystyle \sum_{s}} Q_s)_\ast  = (\hbox{something})_\ast
      (\pi^{A_j}_{\ell})_\ast(\hbox{something}){}_\ast\colon \ \  Z^r(X)\ \to\ Z^r(X)\ .\]      
      
      The right--hand--side of this equality is $0$, since 
         \[ (\pi^{A_j}_{\ell})_\ast \Bigl( {A^\ast_{num}(A_j)\over A^\ast_\otimes(A_j)}\Bigr)\]
    for $\ell$ odd (this follows from theorem \ref{skew}, combined with the fact that the $\pi^{A_j}_\ell$ project to odd gradeds of the Beauville filtration on $A^\ast(A_j)$ \cite{DM}, \cite{Sc}). As we have seen in equality (\ref{power}), the left--hand--side is the identity. We conclude that 
      \[ Z^r(X)=0\ .\]      
   \end{proof}

\section{Examples}

In this section, we aim to give some content to theorem \ref{main}, by providing examples of varieties satisfying the assumptions. For convenience, we will write $X^n_d$ for the Fermat hypersurface of dimension $n$ and degree $d$.

\begin{corollary}\label{ex} Let $X$ be one of the following:

\noindent
(1) a Fermat hypersurface $X^n_d$ with $n$ odd;

\noindent
(2) a product $Y_1\times\cdots\times Y_s\times X^n_d$, where the $Y_i$ are varieties with $A^\ast_{hom}(Y_i)=0$ (examples of such varieties can be found in \cite{BKL}, \cite{PW}, \cite{V8}), and $n$ is odd;

\noindent
(3) a product $Y_1\times\cdots\times Y_s\times Y$, where the $Y_i$ are as in (2), and $Y$ is a Calabi--Yau threefold with motive of abelian type (examples of such $Y$ are given in \cite[Section 2]{des} and in \cite{exCY});

\noindent
(4) a product $S\times X^n_{d}$ where $n$ is odd, and $S$ is a regular surface with motive of abelian type (e.g., $S$ can be $X^2_{d^\prime}$, or a double plane branched along $6$ lines in general position \cite{Par}, or a $K3$ surface with Picard number $\ge 19$, or any of the surfaces in \cite{GaP}, \cite{Bonf});

\noindent
(5) a product $Y\times C$, where $C$ is a curve and $Y=X^4_7/\mu_7$ is the fourfold studied in \cite[Proposition 2.17]{V4};

\noindent
(6) a product $Y\times S$, where $S$ is a surface with $A^2_{AJ}(S)=0$, and $Y=X^4_7/\mu_7$ is the fourfold of \cite[Proposition 2.17]{V4};

\noindent
{(7)} a product $S\times Y$, where $S$ is a regular surface with motive of abelian type, and $Y$ is a Calabi--Yau threefold with motive of abelian type; 


\noindent
{(8)} the Calabi--Yau $5$--fold obtained from a product of $5$ elliptic curves as in \cite[Corollary 2.3]{CH}.

Then 
  \[ A^r_\otimes(X)=A^r_{num}(X)\ \ \hbox{for\ all\ }r\ .\]
\end{corollary}

\begin{proof} Clearly, all these examples have motive of abelian type: for case (1), this follows from Shioda's inductive structure \cite{Shi}; for case (2) this follows from \cite[Theorem 5]{V2} (or, independently, from \cite{Kim2}); case (5) follows from \cite[Proposition 2.17]{V4}; the surfaces in case (6) follow from \cite[Theorem 4]{V2}. It remains to check hypothesis (\rom2) of theorem \ref{main} is verified. For cases (1), (2), (3) this is clear since in these cases the even--degree cohomology is algebraic, and so
  \[ H^{2i}(X)=N^i H^{2i}(X)=\wt{N}^i H^{2i}(X)\subset \wt{N}^{i-1} H^{2i}(X)\ .\]
  
  In case (4), we have
   \[  H^{2i}(X)=\bigoplus_{k+\ell=2i} H^k(S)\otimes H^{\ell}(X^n_{d^\prime})\ .\]
 Any direct summand with $k\not=2$ consists of algebraic classes. For $k=2$, we have
   \[ H^2(X^2_d)\otimes  H^{2i-2}(X^n_{d^\prime})\ \subset\ H^2(X^2_d)\otimes \wt{N}^{i-1} H^{2i-2}(X^n_{d^\prime})\ \subset\   \wt{N}^{i-1} H^{2i}(X^2_d\times X^{n}_{d^\prime})\ .\]  
 
 In case (5), we have $H^4(Y)=\wt{N}^1 H^4(Y)$ \cite[Proposition 2.17]{V4}. It follows that
   \[ H^4(Y\times C)= H^4(Y)\otimes H^0(C)\oplus H^2(Y)\otimes H^2(C)\ \subset\ \wt{N}^1 H^4(Y\times C)\ .\]
   
   In case (6), we have 
     \[ \begin{split} H^4(X)&= H^4(Y)\otimes H^0(S)\oplus H^2(Y)\otimes H^2(S)\oplus H^0(Y)\otimes H^4(S)\\ 
                               &\subset\ \wt{N}^1 H^4(Y)\otimes H^0(S)\oplus \wt{N}^1 H^2(Y)\otimes \wt{N}^1 H^2(S)\oplus H^0(Y)\otimes \wt{N}^2 H^4(S)\\
                               &\subset \wt{N}^1 H^4(X)\ ,\\
                               \end{split}\]
                   and likewise
 \[ \begin{split} H^6(X)&= H^6(Y)\otimes H^0(S)\oplus H^4(Y)\otimes H^2(S)\oplus H^2(Y)\otimes H^4(S)  \\ 
                               &\subset\ \wt{N}^2 H^6(Y)\otimes H^0(S)\oplus \wt{N}^1 H^4(Y)\otimes \wt{N}^1 H^2(S)\oplus \wt{N}^1 H^2(Y)\otimes \wt{N}^2 H^4(S)\\
                               &\subset \wt{N}^2 H^6(X)\ .\\
                               \end{split}\]

 Cases (7) is similar to case (4).
 
As to case (8): let $E_1,\ldots, E_5$ be elliptic curves, and let $X$ be a Calabi--Yau $5$--fold obtained as a smooth model of the quotient
  \[   (E_1\times\cdots\times E_5)/\ZZ_2^4\]
  as in \cite[Corollary 2.3]{CH}. It is readily checked (using the argument of \cite[Lemma 2.4]{CH}) that
  \[  H^4(E_1\times\cdots\times E_5)^{\ZZ_2^4}\subset N^2 H^4(E_1\times\cdots\times E_5)\ . \]
Next, the inductive construction of \cite[Proposition 2.1]{CH}  shows $X$ is of the form $Z/\ZZ_2^4$, where $Z$ is obtained from $E_1\times\cdots\times E_5$ by blowing up some rational subvarieties. Since rational varieties of dimension $\le 3$ verify the Lefschetz standard conjecture, this implies
  \[  H^i(Z)\ \subset\ \ima \bigl( H^i(E_1\times\cdots\times E_5)\ \to\ H^i(Z)\bigr) \ \cup \ \wt{N}^1 H^i(Z)\ \ \ \hbox{for\ all\ }i\ .\]
  In particular, it follows that
  \[ H^4(X)=H^4(Z)^{\ZZ_2^4}\ \subset\   \ \wt{N}^1 H^4(X)\ .\]
     \end{proof}

\vskip1cm

\begin{nonumberingt} The ideas developed in this note grew into being during the Strasbourg 2014---2015 groupe de travail based on the monograph \cite{Vo}. Thanks to all the participants of this groupe de travail for the pleasant and stimulating atmosphere. 
Thanks to Charles Vial and an anonymous referee for helpful comments.
Many thanks to Yasuyo, Kai and Len for not disturbing when the door to my office is closed.
\end{nonumberingt}

\end{document}